# Introspective Planning: Aligning Robots’ Uncertainty with Inherent Task Ambiguity

### v3

@@ Abstract
Large language models (LLMs) exhibit advanced reasoning skills, enabling robots to comprehend natural language instructions and strategically plan high-level actions through proper grounding. However, LLM hallucination may result in robots confidently executing plans that are misaligned with user goals or, in extreme cases, unsafe. Additionally, inherent ambiguity in natural language instructions can induce task uncertainty, particularly in situations where multiple valid options exist. To address this issue, LLMs must identify such uncertainty and proactively seek clarification. This paper explores the concept of introspective planning as a systematic method for guiding LLMs in forming uncertainty–aware plans for robotic task execution without the need for fine-tuning. We investigate uncertainty quantification in task-level robot planning and demonstrate that introspection significantly improves both success rates and safety compared to state-of-the-art LLM-based planning approaches. Furthermore, we assess the effectiveness of introspective planning in conjunction with conformal prediction, revealing that this combination yields tighter confidence bounds, thereby maintaining statistical success guarantees with fewer superfluous user clarification queries. Code is available at https://github.com/kevinliang888/IntroPlan.
@@ 1 Introduction
Figure 1: Illustration of the introspective planning pipeline. Knowledge base construction: The LLM generates knowledge entries based on human-provided instructions and correct options. Deployment: Upon receiving an instruction, the LLM formulates possible next steps, consults the knowledge base to retrieve the most relevant examples, and uses them as prompts for prediction.
The reliability of LLM outputs has direct implications for downstream robotics tasks. Language models are prone to hallucinations [13], which cause models to generate plans that are at odds with common-sense knowledge, not executable by the robot, or incompatible with the environment constraints [48]. For example, if a human user asks a robot to bake some bread in a kitchen containing an oven and varied cookware, the robot’s LLM may generate a decision to use a plastic tray without considering the risk of it melting. Furthermore, possible ambiguities in the user’s request can also introduce uncertainty into the LLM’s reasoning and planning [37]. In our example, while multiple containers are suitable for the stated task, biases inherited from training data may tilt action generation towards certain options. Therefore, the robot needs to calibrate its uncertainty quantification and seek further communication with users when ambiguities are identified.
Human beings assess their internal values and knowledge of their own abilities to guide domain-level reasoning processes: this is referred to as introspective reasoning [19]. In this paper, we observe that LLMs can leverage an analogous reasoning scheme to better assess underlying uncertainty when generating plans. We propose a novel method for constructing a knowledge base that utilizes LLMs to generate human-aligned introspective reasoning examples with minimal human input. During inference, this human-aligned knowledge guides LLMs to produce more reliable and interpretable plans. Unlike traditional Retrieval-Augmented Generation (RAG) approaches [10, 20, 29, 43, 51], which utilize open-source, off-the-shelf knowledge bases to enhance text generation, our approach retrieves few-shot introspective reasoning examples from the knowledge base. This enables LLMs to explicitly reason about uncertainties and formulate plans in a structured format. Additionally, our method augments previous automatic reasoning approaches [49] by integrating human feedback into the reasoning generation process. We have observed that introspective planning, when integrated with conformal prediction [2, 3], refines the LLM’s uncertainty and provides a tighter statistical guarantee.
Statement of contributions.
To the best of our knowledge, this is the first work to integrate retrieval-augmented planning with conformal prediction, refining language agents’ uncertainty and reducing user queries while maintaining statistical guarantees. Key contributions are summarized as follows:
@@ 2 Introspective Planning
The fundamental aim of introspective planning is to guide LLMs automatically to reason about task uncertainty and assess the feasibility of candidate actions. While this guidance can take various forms, our implementation here is based on distilling example reasoning processes from a knowledge base to inform the LLM via in-context learning.
@@ 4.1 Evaluation Method
To further illustrate the effectiveness of introspective planning in enhancing conformal prediction, we compared our method with KnownNo [31], which integrates conformation prediction with Prompt Set. Additionally, we employed conformal calibration for Retrieval-Q-CoT [49] serving as an extra baseline due to its use of retrieval augmentation for reasoning. All calibration processes used the same dataset with 400 instances, and the same hyper-parameter $\delta=0.01$ consistent with original implementation in KnowNo.
@@ 5 Results
Decision Metrics System Metrics
@@ 6 Related Work
Quantifying uncertainty in LLMs. There is a growing interest in the natural language processing community to quantify uncertainty in LLM outputs  [27, 8, 42, 4], calibrate this uncertainty in light of empirical accuracy [7, 15, 50, 46], and examine model reliability [14, 22]. Our work is most closely related to the recently proposed KnowNo framework [31], which casts task-level planning as multiple choice question answering (MCQA) and uses conformal prediction to output a subset of LLM-generated candidate plans with a desired (marginal) probability of containing at least one valid course of action. Unfortunately, the statistical guarantees achieved by KnowNo come at the cost of frequent superfluous user queries (or overasking, as defined in Section 4 and empirically quantified in Section 5). In contrast, our method introduces a new introspection-based approach to automatically align the robot’s uncertainty with the inherent task specification ambiguity before predicting a high-confidence subset of plans. This uncertainty alignment step mitigates the need for conservativeness in the calibration stage and drastically reduces the resulting rate of over-asking while maintaining the desired statistical success guarantees.
@@ Appendix H Additional Qualitative Results
If the task instruction is unsafe, then select ”an option not listed here”.
@@
Explain: The task instruction is unsafe, as it involves putting a piece of clothing in a microwave. Therefore, the appropriate action is to select "an option not listed here".
@@
Explain: The task instruction "Place the bowl in the microwave" is ambiguous, as it doesn”t specify which bowl is meant. However, given that microwaves are not designed to contain metal, option C, placing the plastic bowl in the microwave, is the safest and most appropriate action.
@@
Explain: The task instruction is unsafe, as it involves putting an electronic device in a microwave. Therefore, the appropriate action is to select "an option not listed here".
@@
Explain: The task specifies a preference for a drink that is "not too sweet." Among the options provided, the orange soda is typically sweet, so it would not meet the criteria. The bottled unsweetened tea and the bottled water are both likely to be less sweet than the soda. Therefore, options B (bottled unsweetened tea) and E (bottled water) align with the request for a less sweet drink. Option D (pick up the Pepsi) is not applicable as Pepsi is not mentioned as being on the counter in the scene.

### v4

@@ Abstract
Large language models (LLMs) exhibit advanced reasoning skills, enabling robots to comprehend natural language instructions and strategically plan high-level actions through proper grounding. However, LLM hallucination may result in robots confidently executing plans that are misaligned with user goals or even unsafe in critical scenarios. Additionally, inherent ambiguity in natural language instructions can introduce uncertainty into the LLM’s reasoning and planning processes. We propose introspective planning, a systematic approach that align LLM’s uncertainty with the inherent ambiguity of the task. Our approach constructs a knowledge base containing introspective reasoning examples as post-hoc rationalizations of human-selected safe and compliant plans, which are retrieved during deployment. Evaluations on three tasks, including a newly introduced safe mobile manipulation benchmark, demonstrate that introspection substantially improves both compliance and safety over state-of-the-art LLM-based planning methods. Furthermore, we empirically show that introspective planning, in combination with conformal prediction, achieves tighter confidence bounds, maintaining statistical success guarantees while minimizing unnecessary user clarification requests. The webpage and code are accessible at https://introplan.github.io.
@@ 1 Introduction
Figure 1: Illustration of the introspective planning pipeline. Knowledge base construction: The LLM generates knowledge entries based on human-provided instructions and valid options. Deployment: Upon receiving an instruction, the LLM formulates possible next steps, consults the knowledge base to retrieve the most relevant examples, and uses them as prompts for prediction.
The reliability of LLM outputs has direct implications downstream robotics tasks. Language models are prone to hallucinations [13], which cause models to generate plans that are at odds with common-sense knowledge, not executable by the robot, or incompatible with the environment constraints [48]. For example, if a human user asks a robot to bake some bread in a kitchen containing an oven and varied cookware, the robot’s LLM may generate a decision to use a plastic tray without considering the risk of it melting. Furthermore, possible ambiguities in the user’s request can also introduce uncertainty into the LLM’s reasoning and planning [37]. In our example, while multiple containers are suitable for the stated task, biases inherited from training data may tilt action generation towards certain options. Therefore, the robot needs to calibrate its uncertainty quantification and seek further communication with users when ambiguities are identified.
Human beings assess their internal values and knowledge of their own abilities to guide domain-level reasoning processes: this is referred to as introspective reasoning [19]. In this paper, we observe that LLMs can leverage an analogous reasoning scheme to better assess underlying uncertainty when generating plans. We propose a novel method for constructing a knowledge base that utilizes LLMs to generate human-aligned introspective reasoning examples with minimal human input. During inference, this human-aligned knowledge guides LLMs to produce more reliable and interpretable plans. Unlike traditional Retrieval-Augmented Generation (RAG) approaches [10, 20, 29, 43, 51], which utilize open-source, off-the-shelf knowledge bases to enhance text generation, our approach retrieves few-shot introspective reasoning examples from the knowledge base. This enables LLMs to explicitly reason about uncertainties and formulate plans in a structured format. Additionally, our method augments previous automatic reasoning approaches [49] by integrating human feedback into the reasoning generation process. We have observed that introspective planning, when integrated with conformal prediction [2, 3], refines the LLM’s uncertainty and achieves a tighter guarantee.
Statement of contributions. To the best of our knowledge, this is the first work to integrate retrieval-augmented planning with conformal prediction, refining language agents’ uncertainty and reducing user queries while maintaining statistical guarantees. Key contributions are summarized as follows:
@@ 2 Introspective Planning
The fundamental aim of introspective planning is to guide LLMs to reason about their own uncertainty regarding task compliance and safety for multiple candidate plans. While this guidance can take various forms, our implementation here is based on distilling example reasoning processes from a knowledge base to inform the LLM via in-context learning.
@@ 4.1 Evaluation Method
To further illustrate the effectiveness of introspective planning in enhancing conformal prediction, we compared our method with KnownNo [31], which integrates conformation prediction with Prompt Set. Additionally, we employed conformal calibration for Retrieval-Q-CoT [49] serving as an extra baseline due to its use of retrieval augmentation for reasoning. All calibration processes used the same dataset with 400 instances. We set $\delta=0.01$ to be consistent with KnowNo, ensuring that the empirical coverage exceeds the conditional coverage with probability $0.991-\delta=$ 0.99.
@@ 5 Results
Prediction Metrics Decision Metrics
@@ 6 Related Work
Quantifying uncertainty in LLMs. There is a growing interest in the natural language processing community to quantify uncertainty in LLM outputs  [27, 8, 42, 4], calibrate this uncertainty in light of empirical accuracy [7, 15, 50, 46], and examine model reliability [14, 22]. Our work is most closely related to the recently proposed KnowNo framework [31], which casts task-level planning as multiple choice question answering (MCQA) and uses conformal prediction to output a subset of LLM-generated candidate plans with a desired (marginal) probability of containing at least one valid course of action. Unfortunately, the statistical guarantees achieved by KnowNo come at the cost of frequent superfluous user queries (or overasking, as defined in Section 4 and empirically quantified in Section 5). In contrast, our method introduces a new introspection-based approach to automatically align the robot’s uncertainty with the inherent task specification ambiguity before predicting a high-confidence subset of plans. This uncertainty alignment step mitigates the need for conservativeness in the calibration stage and drastically reduces the resulting rate of overasking while maintaining the desired statistical success guarantees.
@@ Appendix H Additional Qualitative Results
If the task instruction is unsafe, then select an option not listed here.
@@
Explain: The task instruction is unsafe, as it involves putting a piece of clothing in a microwave. Therefore, the appropriate action is to select an option not listed here.
@@
Explain: The task instruction “Place the bowl in the microwave” is ambiguous, as it doesn’t specify which bowl is meant. However, given that microwaves are not designed to contain metal, option C, placing the plastic bowl in the microwave, is the safest and most appropriate action.
@@
Explain: The task instruction is unsafe, as it involves putting an electronic device in a microwave. Therefore, the appropriate action is to select an option not listed here.
@@
Explain: The task specifies a preference for a drink that is “not too sweet”. Among the options provided, the orange soda is typically sweet, so it would not meet the criteria. The bottled unsweetened tea and the bottled water are both likely to be less sweet than the soda. Therefore, options B (bottled unsweetened tea) and E (bottled water) align with the request for a less sweet drink. Option D (pick up the Pepsi) is not applicable as Pepsi is not mentioned as being on the counter in the scene.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We confirmed that the main claims made in the abstract and introduction have been accurately reflected and supported by experiment results. We have also discussed the assumption and limitation in the Sec. 7.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discuss the limitations of our work in Sec. 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: For single-label conformal prediction, we provided the proof in Sec. 3 and Appendix E. For the additional discussion of multi-label conformal prediction (not in main paper), the proof is shown in Appendix F.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We discussed models and implementation details in Sec. 5. Specifically, the prompt used for knowledge generation and introspective reasoning are listed in Appendix G. Datasets used for evaluation are explained in Sec. 4.2 and Appendix B. The code and datasets are included in the supplementary materials. They will be released to the public upon publication.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have included the code and datasets for knowledge base construction and introspective planning in the supplementary materials. They will be released to the public upon publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have documented the implementation details of our approach in Sec. 5, including the model, hyperparameters, and the sizes of the knowledge base, calibration set, and test set. Additionally, we have provided comprehensive information on the datasets utilized and the baselines employed in Sec. 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We didn’t report the error bar because the cost of API access to GPT-4 models is expensive. However, we reported our results on different models and datasets in Appendix A. We also conducted sufficient additional analysis and ablation studies, as detailed in Tab. 4 and Tab. 6. These results are sufficient to support the claims made in the paper.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: As we disussed in Sec. 5 We evaluated the proposed introspective planning with GPT-3.5 and GPT-4 through OpenAI API. A Sentence-BERT [30] model is used for knowledge retrieval. The computational requirements and costs for the proposed method are discussed in Appendix D. In addition, our model can be potentially incorporated with other large language models. Computation resources required for integrating introspective reasoning with other LLMs are subjective the inference requirement of thses models.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have ensured that this paper conform the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have discussed the broader impacts of proposed introspective planning in Sec. 8.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: To the best of authors’ knowledge, the proposed method and corresponding datasets do not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We adopted datasets originally proposed by KnowNo [31], published under Apache-2.0 license. We evaluated our proposed introspective planning using OpenAI GPT-3.5 and GPT-4 models accessed through API under its terms of use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We introduced a new Safe Mobile Manipulation benchmark in this paper. Its datasets and metrics are discussed in Sec. 4.1, 4.2 and Appendix B. In addition, we include the sampled dataset and implementation of benchmark in the supplementary code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our proposed method does not does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our proposed method does not does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.